OpenStudy (anonymous):

Which of the following are arithmetic sequences? a. -2,2,6,10 b. 1,3,9,27 c. 5,10,20,40 d. 5,1,-3,-7

OpenStudy (amistre64):

how do we define a sequence as arithmetic?

OpenStudy (texaschic101):

In arithmatic sequences, the difference is found by addition. In geometric sequences, the differences are found by multiplication

OpenStudy (anonymous):

If a sequence of values follows a pattern of adding a fixed amount from one term to the next, it is referred to as an arithmetic sequence

OpenStudy (anonymous):

Please help

OpenStudy (amistre64):

take the terms and see if you add the same amount to the first one to get the to second one, as you do to get from the second to third

OpenStudy (amistre64):

\[t_2-t_1=n\] \[t_3-t_2=k\] if n=k, your good to go in general

OpenStudy (texaschic101):

look at your answer choices you have to find the pattern. can you add anything to the first term to get to the second term ? can you add the same number to the second term to get to the third term ?

OpenStudy (amistre64):

of course order is immaterial as long as you get the same value ..... for example: -2,2,6,10 -2 - 2 = -4 2 - 6 = -4 6 - 10 = -4 that is one of them ... i believe there is one more

OpenStudy (amistre64):

1,3,9,27 1 - 3 = -2 3 - 9 = -6 so this isnt the same each time ... try the others to test them out

OpenStudy (texaschic101):

Here is an example of an arithmetic sequence... 3,5,7,9,11 3 + 2 = 5 5 + 2 = 7 7 + 2 = 9 your common difference is 2 (by addition) geometric sequence example 3,6,12,24 3 x 2 = 6 6 x 2 = 12 12 x 2 = 24 your common difference is 2 (by multiplication)

OpenStudy (anonymous):

i might sound really stupid but how exactly do you do this

OpenStudy (anonymous):

ok so you added two to get from 3 to 5 and 5 to seven so its added by two ok i get that

OpenStudy (texaschic101):

look at your first answer choice... -2,2,6,10 take your second term and subtract your first term -2 - (-2) = 4 now take your third term and subtract your second term 6 - 2 = 4 4th term subtract 3rd term 10 - 6 = 4 common difference is 4, and you can add 4 to each term to get the next number. This is an arithmetic sequence.

OpenStudy (anonymous):

okay i get it know thankyou

OpenStudy (anonymous):

now

OpenStudy (texaschic101):

glad to help :)